An analyst has the following information about the Highly Smart REIT valuation:

  • Estimated NOI next year                                                              $65,000
  • Share price $33.00
  • Total debt & liabilities $750,000
  • Net income $45,000
  • Depreciation $5,000
  • Profits on sale of Property                                                            $6,000
  • Recurring capital expenditures $8,000

Other information

  • Share outstanding 10,000
  • Earnings per share $4.50
  • Average industry P/FFO multiple 8x
  • Average industry P/AFFO multiple 10x

Based on the given information, estimate the value of the Highly Smart REIT using:

  1. P/FFO and
  2. P/AFFO

(Note: FFO = Net Income + property depreciation – property sales profits

AFFO = FFO – recurring capital expenditures)

Evaluate whether Highly Smart REIT is overvalued or undervalued? (20 Marks)
Answer #1

1.

FFO = Net Income + property depreciation – property sales profits

FFO = $45000 + 5000 - 6000

FFO = $44000

FFO per share = 44000 / 10000 = $4.4

P/FFO = Price / FFO = $33 / $4.4 = 7.50

2. AFFO = FFO – recurring capital expenditures

AFFO = $44000 - 8000 = $36000

AFFO per share = 36000 / 10000 = $3.60

P/AFFO = Price / AFFO = $33 / $3.60 = 9.17

As the P/FFO ratio of REIT 7.50x is less than industry average of 8X. the company is undervalued.